# Can LLMs Convert Graphs to Text-Attributed Graphs?

### v1

@@ Abstract
Graphs are ubiquitous data structures found in numerous real-world applications, such as drug discovery, recommender systems, and social network analysis. Graph neural networks (GNNs) have become a popular tool to learn node embeddings through message passing on these structures. However, a significant challenge arises when applying GNNs to multiple graphs with different feature spaces, as existing GNN architectures are not designed for cross-graph feature alignment. To address this, recent approaches introduce text-attributed graphs, where each node is associated with a textual description, enabling the use of a shared textual encoder to project nodes from different graphs into a unified feature space. While promising, this method relies heavily on the availability of text-attributed data, which can be difficult to obtain in practice. To bridge this gap, we propose a novel method named Topology-Aware Node description Synthesis (TANS), which leverages large language models (LLMs) to automatically convert existing graphs into text-attributed graphs. The key idea is to integrate topological information with each node’s properties, enhancing the LLMs’ ability to explain how graph topology influences node semantics. We evaluate our TANS on text-rich, text-limited, and text-free graphs, demonstrating that it enables a single GNN to operate across diverse graphs. Notably, on text-free graphs, our method significantly outperforms existing approaches that manually design node features, showcasing the potential of LLMs for preprocessing graph-structured data, even in the absence of textual information. The code and data are available at https://github.com/Zehong-Wang/TANS.
@@ 1 Introduction
Graph-structured data are prevalent in many real-world domains, including chemistry, social networks, and recommendation systems (Gilmer et al., 2017; Wang et al., 2023; Zhao et al., 2023; Zhang et al., 2024). Graph neural networks (GNNs) (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Gilmer et al., 2017) have emerged as powerful tools for processing such data by learning node representations that capture both the structural and feature properties of graphs. In each GNN layer, the model first applies a linear transformation to project the input node features or previous layer embeddings into a new space. Then, through message passing (Gilmer et al., 2017), the model aggregates information from neighboring nodes to update node embeddings.
@@
Although GNNs have shown remarkable success across various applications, a significant limitation is their inability to handle multiple graphs jointly, especially when these graphs have different node feature dimensions (Liu et al., 2024a; Wang et al., 2024c), as illustrated in Figure 1 (a). This is primarily because the input dimension of message-passing GNNs is fixed, making it challenging to accommodate graphs with varying feature spaces. In practice, many real-world graphs possess diverse feature dimensions, which prevents a single GNN from being applied across multiple graphs. This limitation poses a substantial barrier to deploying GNNs in scenarios requiring cross-graph learning, such as transfer learning (Han et al., 2021; Zhu et al., 2021; Wang et al., 2024d), domain adaptation (Dai et al., 2022; Mao et al., 2024), out-of-distribution detection (Li et al., 2022), and the development of graph foundation models (Liu et al., 2024a; Wang et al., 2024c). Overcoming this challenge is crucial for extending the applicability of GNNs to more complex and diverse real-world tasks.
@@
To address the challenge of aligning node features across multiple graphs, two main approaches have been proposed. The first approach involves using singular value decomposition (SVD) (Yu et al., 2024; Zhao et al., 2024) to decompose the original node features of different graphs, aligning the number of dimensions across graphs. However, this method has two limitations: (1) SVD is only applicable to graphs with node features, making it ineffective for graphs that lack such features. (2) While SVD aligns the feature dimensions, it does not ensure that the semantic meaning of the features is consistent across different graphs (Yu et al., 2024). A different approach introduces the concept of text-attributed graphs (TAGs) (Yan et al., 2023), where each node is associated with a textual description. Researchers use textual encoders (Liu et al., 2024a; Wang et al., 2024c) to transform these descriptions into textual embeddings, better aligning node features across graphs and enabling a single GNN to handle multiple graphs, as shown in Figure 1 (b). Compared to SVD, this method provides more meaningful feature alignment and can generalize to unseen TAGs by simply processing the node descriptions (Wang et al., 2024c). However, despite these advantages, TAGs face a significant limitation: the collection of high-quality textual descriptions for all nodes is often impractical, making it challenging to apply this method in real-world scenarios (Chen et al., 2024b).
@@
In this paper, we address the challenge of collecting textual descriptions for feature alignment in TAGs. Inspired by the success of LLMs in generating synthetic data (Tang et al., 2023; Long et al., 2024), we propose a novel method named Topology-Aware Node description Synthesis (TANS) that leverages LLMs to automatically generate node-level textual descriptions for existing graph datasets, acting as the first work towards this direction. TANS focuses on the node classification task and uses the inherent topological information (Bondy and Murty, 2008) of each node to guide the LLMs in producing meaningful textual descriptions that capture the role of each node within the graph. Our method is versatile, as it not only applies to graphs without textual descriptions but also enhances graphs with existing rich or limited textual data by generating more informative descriptions. We evaluate our approach on five diverse graph datasets, covering text-rich, text-limited, and text-free scenarios, each with distinct scales and topologies. The experimental results demonstrate the effectiveness of our method, showing superior performance in generating node properties. Moreover, our approach is robust in transfer learning and domain adaptation tasks, further showcasing the value of LLM-generated descriptions in aligning node features across different graphs.
Figure 2: The framework of our proposed TANS.
@@ 2 Backgrounds
${\bm{z}}_{i}^{(l)}=\sigma\Bigl{(}{\bm{W}}_{1}{\bm{z}}_{i}^{(l-1)}+{\bm{W}}_{2}\Bigl{(}\frac{1}{|{\mathcal{N}}(i)|}\sum_{j\in{\mathcal{N}}(i)}{\bm{z}}_{j}^{(l-1)}\Bigr{)}\Bigr{)},$
@@ Appendix A More Related Works
Graph Neural Networks. Graph neural networks (GNNs) (Liu et al., 2023; Wang et al., 2024b, e; Liu et al., 2024b) are effective in various graph learning tasks by utilizing the message passing framework. For example, GCN (Kipf and Welling, 2017) leverages the Laplacian matrix for message passing, MPNN (Gilmer et al., 2017) formally defines the message passing framework, GraphSAGE (Hamilton et al., 2017) extends it to inductive learning, and GAT (Veličković et al., 2018) introduces attention mechanisms. Further works (Wang et al., 2019, 2023; Zhang et al., 2024) have extended message passing to various graph types and applications. However, a key limitation of message passing GNNs is their inability to handle graphs with different feature spaces (Liu et al., 2024a; Wang et al., 2024c), highlighting the need for effective feature alignment methods across graphs.

### v2

@@ Abstract
Graphs are ubiquitous structures found in numerous real-world applications, such as drug discovery, recommender systems, and social network analysis. To model graph-structured data, graph neural networks (GNNs) have become a popular tool. However, existing GNN architectures encounter challenges in cross-graph learning where multiple graphs have different feature spaces. To address this, recent approaches introduce text-attributed graphs (TAGs), where each node is associated with a textual description, which can be projected into a unified feature space using textual encoders. While promising, this method relies heavily on the availability of text-attributed graph data, which is difficult to obtain in practice. To bridge this gap, we propose a novel method named Topology-Aware Node description Synthesis (TANS), leveraging large language models (LLMs) to convert existing graphs into text-attributed graphs. The key idea is to integrate topological information into LLMs to explain how graph topology influences node semantics. We evaluate our TANS on text-rich, text-limited, and text-free graphs, demonstrating its applicability. Notably, on text-free graphs, our method significantly outperforms existing approaches that manually design node features, showcasing the potential of LLMs for preprocessing graph-structured data in the absence of textual information. The code and data are available at https://github.com/Zehong-Wang/TANS.
@@ 1 Introduction
Graph-structured data are prevalent in many real-world domains, including chemistry (Zhao et al., 2023), social networks (Zhang et al., 2024a), and recommendation systems (Zhang et al., 2024b). Graph neural networks (GNNs) (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Gilmer et al., 2017; Wang et al., 2023) have emerged as powerful tools for processing such data by learning node representations that capture both the structural and attributed properties of graphs. In each GNN layer, the model first applies a linear transformation to project the input node features or previous layer embeddings into a new space. Then, through message passing (Gilmer et al., 2017), the model aggregates information from neighboring nodes to update node embeddings.
@@
Although GNNs have shown remarkable success across various applications, a significant limitation is their inability to handle multiple graphs jointly, especially when these graphs have different node feature dimensions (Liu et al., 2024a; Wang et al., 2024c), as illustrated in Figure 1 (a). This is primarily because the input dimension of message-passing GNNs is fixed, making it challenging to accommodate graphs with varying feature spaces. In practice, many real-world graphs possess diverse feature dimensions, preventing a single GNN from being applied across multiple graphs. This limitation poses a substantial barrier to deploying GNNs in scenarios requiring cross-graph learning, such as transfer learning (Wang et al., 2024d), domain adaptation (Dai et al., 2022), out-of-distribution detection (Li et al., 2022), and graph foundation models (Wang et al., 2024c). Overcoming this challenge is crucial for extending the applicability of GNNs to more complex real-world tasks.
@@
To address the challenge of aligning node features across multiple graphs, two main approaches have been proposed. The first approach involves using singular value decomposition (SVD) (Yu et al., 2024; Zhao et al., 2024) to decompose the original node features of different graphs, aligning the number of dimensions across graphs. However, this method has two limitations: (1) SVD is only applicable to graphs with node features, making it ineffective for graphs that lack such features. (2) While SVD aligns the feature dimensions, it does not ensure that the semantic meaning of the features is consistent across different graphs (Yu et al., 2024). A different approach introduces the concept of text-attributed graphs (TAGs) (Yan et al., 2023), where each node is associated with a textual description. Researchers use textual encoders (Wang et al., 2024c) to transform these descriptions into textual embeddings, better aligning node features across graphs and enabling a single GNN to handle multiple graphs, as shown in Figure 1 (b). Compared to SVD, this method provides more meaningful feature alignment and can generalize to unseen TAGs by simply processing the node descriptions (Wang et al., 2024c). However, despite these advantages, TAGs face a significant limitation: the collection of high-quality textual descriptions for all nodes is often impractical, making it challenging to apply this method in real-world scenarios.
@@
In this paper, we address the challenge of collecting textual descriptions for feature alignment in TAGs. Inspired by the success of LLMs in generating synthetic data (Tang et al., 2023; Long et al., 2024), we propose a novel method named Topology-Aware Node description Synthesis (TANS) that leverages LLMs to automatically generate node-level textual descriptions for existing graph datasets (Tan et al., 2024), acting as the first work towards this direction. TANS focuses on the node classification task and uses the inherent topological information (Bondy and Murty, 2008) of each node to guide the LLMs in producing meaningful textual descriptions that capture the role of each node within the graph. Our method is versatile, as it not only applies to graphs without textual descriptions but also enhances graphs with existing rich or limited textual data by generating more informative descriptions. We evaluate our approach on five diverse graph datasets, covering text-rich, text-limited, and text-free scenarios, each with distinct scales and topologies. The experimental results demonstrate the effectiveness of our method, showing superior performance in generating node properties. Moreover, our approach is robust in transfer learning and domain adaptation, further showing the value of LLM-generated descriptions in aligning node features across different graphs.
Figure 2: The framework of our topology-aware node description synthesis (TANS).
@@ 2 Backgrounds
${\bm{z}}_{i}^{(l)}=\sigma\Bigl{(}{\bm{W}}_{1}{\bm{z}}_{i}^{(l-1)}+\frac{{\bm{W}}_{2}}{|{\mathcal{N}}(i)|}\Bigl{(}\sum_{j\in{\mathcal{N}}(i)}{\bm{z}}_{j}^{(l-1)}\Bigr{)}\Bigr{)},$
@@ 6 Conclusion
Acknowledgement
This work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2340346, IIS-2217239, CNS-2426514, CNS-2203261, and CMMI-2146076. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix A More Related Works
Graph Neural Networks. Graph neural networks (GNNs) (Liu et al., 2023; Wang et al., 2024b, e; Liu et al., 2024b) are effective in various graph learning tasks by utilizing the message passing framework. For example, GCN (Kipf and Welling, 2017) leverages the Laplacian matrix for message passing, MPNN (Gilmer et al., 2017) formally defines the message passing framework, GraphSAGE (Hamilton et al., 2017) extends it to inductive learning, and GAT (Veličković et al., 2018) introduces attention mechanisms. Further works (Wang et al., 2019, 2023; Zhang et al., 2024a) have extended message passing to various graph types and applications. However, a key limitation of message passing GNNs is their inability to handle graphs with different feature spaces (Liu et al., 2024a; Wang et al., 2024c), highlighting the need for effective feature alignment methods across graphs.